A researcher claims that the average salary of a private school teacher is greater than P35,000 with


a standard deviation of P7,000. A sample of 35 teachers has a mean salary of P37,000. Test the


claim of the researcher. At 0.05 level of significance, test the claim of the researcher.

1
Expert's answer
2022-04-21T10:17:47-0400

"H_0:\mu=35000."

"H_a:\mu>35000."

Test statistic: "z=\frac{37000-35000}{\frac{7000}{\sqrt{35}}}=1.69."

P-value: "p=P(Z>1.69)=0.0455."

Since the p-value is less than 0.05, reject the null hypothesis and support the claim.